OpenStudy (anonymous):

Use the arc length formula and the given information to find s. Show your work for full credit. r = 20 ft θ = 19° s = ?

OpenStudy (anonymous):

\[\theta=19=\frac{ \pi }{180 }\times19=\frac{ 19\pi }{ 180 }~radians.\] \[\theta=\frac{ s }{ r },s=r \theta=20\times \frac{ 19\pi }{180 }\] \[=\frac{ 19 \pi }{9 }ft\]

OpenStudy (anonymous):

So then would it be 19*3.14 /9 = 59366/9 = 6.63 ft

OpenStudy (triciaal):

ratio angle/360 = arc length/2*pi*r angle =19, r = 20, arc length = s solve for s

OpenStudy (anonymous):

correct.

OpenStudy (anonymous):



OpenStudy (anonymous):

is this what you mean?

OpenStudy (anonymous):

@surjithayer Thanks!

OpenStudy (triciaal):

yes and pi is usually 3.142

OpenStudy (anonymous):

yw